OpenStudy (anonymous):

Write the equation of a line in point-slope form that is parallel to the line y = 3x +6 and passes through the point (-5, 8)

OpenStudy (anonymous):

m=3 since it is parallel to y=3x+6 (parallel is same slope) Passes through (-5,8) so x1=-5 and y1=8. Plug m, x1, and y1 into the point slope form: y-y1=m(x-x1) to get y-8=3(x+5)